X2uxx +2xyuxy+uyy=u


1
Expert's answer
2021-05-07T14:22:25-0400

(Note - nothing is asked in question , so I am giving basic idea to approach these types of problems)


Solution:-

"x^2u_{xx} +2xyu_{xy}+u_{yy}=u"


"On \ solving \ we \ got \\"


"A= x^2 \\
B = 2xy \\
C = 1"


Now use

"b^2-4ac =0"

To check it is which type of curve this is


"4x^2y^2 - 4x^2= 0 \\
4x^2(y^2-1)=0"


"\implies"

"x=0 , y= +1,-1"



After deciding the curve

Also we can find the colonial form.

As several things .